OpenStudy (anonymous):

solve the differential equation: xy^2y'=x+7

OpenStudy (p0sitr0n):

xy^(2y')=x+7 2y'lnxy=ln(x+7) y' = ln(x+7) / 2lnxy ?

OpenStudy (anonymous):

The first part of the equation is x(y'')(y') not xy^(2y')

OpenStudy (p0sitr0n):

you need two equations then i guess ?

OpenStudy (anonymous):

How do i do that?